Computation ionization constant of hydrofluoric acid, 0.1 F solution is 7.13 per cent ionized.


1
Expert's answer
2021-03-15T09:06:35-0400

Hydrofluoric acid dissociates as following:

HF + H2O = H3O+ + F-

Ionization constant equals:

Ka = [H3O+] [F-] / [HF]

As 7.13 % of solution is ionized:

[F-] = [H3O+] = 7.13 % × [HF]initial / 100 % = 0.00713 F

[HF] = 0.1 - 0.00713 = 0.09287 F

Finally:

Ka = 0.00713 F × 0.00713 F / 0.09287 F = 5.47 × 10-4


Answer: 5.47 × 10-4